Hockey teams receive 2 points if they win and 1 point for a tie. One season a team won a championship with 54 points. They won 9 more games than they tied. How many wins and how many ties did the team have?
Help me please I am terrible with these word/math problems.

Answer by scott8148(6628)

let x=ties, so x+9=wins __ 1(x)+2(x+9)=54 __ x+2x+18=54 __ 3x=36 __ x=12

12 ties and 21 wins